OpenStudy (anonymous):

what could be the value of of (g-3k) in the equation (g-3k) (3g-2k)=0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

zero-product-rule if a product of many factors equals 0 then one of them is 0

OpenStudy (anonymous):

ah. okay. thank you :)

OpenStudy (anonymous):

yw,by the way (3g-2k) could also be 0

OpenStudy (anonymous):

okay ^_^

OpenStudy (anonymous):

(g-3k) (3g-2k)=0 is the same as (g-3k)=0 and (3g-2k)=0 :)